# Source-Aware Training Enables Knowledge Attribution in Language Models

### v2

@@ Abstract
Large language models (LLMs) learn a vast amount of knowledge during pretraining, but they are often oblivious to the source(s) of such knowledge. We investigate the problem of intrinsic source citation, where LLMs are required to cite the pretraining source supporting a generated response. Intrinsic source citation can enhance LLM transparency, interpretability, and verifiability. To give LLMs such ability, we explore source-aware training—a post pretraining recipe that involves (i) training the LLM to associate unique source document identifiers with the knowledge in each document, followed by (ii) an instruction-tuning to teach the LLM to cite a supporting pretraining source when prompted. Source-aware training can easily be applied to pretrained LLMs off the shelf, and diverges minimally from existing pretraining/fine-tuning frameworks. Through experiments on carefully curated data, we demonstrate that our training recipe can enable faithful attribution to the pretraining data without a substantial impact on the model’s quality compared to standard pretraining. Our results also highlight the importance of data augmentation in achieving attribution. [^1] Code and data available here: https://github.com/mukhal/intrinsic-source-citation.
@@ 1 Introduction
We focus on intrinsic source citation, where the LLM should cite source documents from the pretraining data from which it acquired its relevant parametric knowledge. Compared to retrieval-based approaches such as RAG (Lewis et al., 2020; Guu et al., 2020) or post-hoc techniques (He et al., 2023; Gao et al., 2023a), intrinsic source citation is inherently tied to the model itself, enables more faithful attribution to its parametric knowledge, thus opens up unique opportunities for improved interpretability (Alvarez Melis & Jaakkola, 2018; Marasovic et al., 2022).
To this end, we explore source-aware training—a post-pretraining recipe that enables a LLM to cite its pretraining data based on its parametric knowledge. Our motivation is three-fold. First, a significant portion of an LLM’s knowledge is acquired during pretraining, therefore citing evidence for this parametric knowledge can greatly enhance the LLM trustworthiness. Second, the standard practice for LLM pretraining neglects the attribution angle, which explains why the current generation of LLMs fails to provide reliable citations (Agrawal et al., 2023; Zuccon et al., 2023). We aim to explore a training procedure that naturally facilitates citation of the pretraining data. Finally, from a scientific perspective, it is intriguing to investigate whether and how current language models can be trained to reference their pretraining data.
We inquire: Given an off-the-shelf LLM, can we train it to attribute its generations to the supporting sources from the pretraining data? Our goal is to cite the pretraining documents themselves (see Figure 1). Our setup mirrors existing frameworks for LLM pretraining and can be summarized as follows: We take an off-the-shelf LLM, continue pretraining it on a corpus associating each document with a unique identifier, then fine-tune it to answer questions about the acquired knowledge while providing citations. The citation is achieved by generating an identifier of a document supporting the answer. Continual pretraining is done as in prior work, with the main difference of injecting the document identifiers into the pretraining data—minimal changes in the model’s architecture or implementation are needed.
To study the generalization over this task and simulate a realistic fine-tuning setting, we limit our instruction tuning stage to a subset of the pretraining documents (in-domain) and evaluate the model’s attribution ability over the remaining (out-of-domain) documents. We run experiments over a synthetic pretraining corpus of fake biographies and show that LLMs can achieve reasonable attribution when answering a question about the out-of-domain documents.
@@
  • We explore a source-aware training recipe that can be applied to off-the-shelf LLMs to give them the ability to attribute their outputs to the pretraining sources. On synthetic data, we show that such training can achieve reasonable attribution while maintaining a good balance with the LLM quality compared to standard pretraining.
  • We examine the impact of various training strategies on attribution such data augmentation, and our findings can inform future efforts to train attribution-capable models at a large scale.
@@ 2 Related Work
Model-based techniques involve prompting the model directly to generate citations for its parametric knowledge (Weller et al., 2023; Zuccon et al., 2023) or scaling techniques such as influence functions (Koh & Liang, 2017) to large models (Grosse et al., 2023). Model-based attribution is arguably more faithful than retrieval-based approaches as the citation mechanism is intrinsic to the model (Alvarez Melis & Jaakkola, 2018; Marasovic et al., 2022). However, standard approaches to pretraining LMs do not take into account the need for the language model to cite its pretraining data, which is where our work comes into play. Bohnet et al. (2022) proposed the task of attributed question-answering and evaluated the attribution performance of different systems using the AutoAIS metric (Rashkin et al., 2023; Gao et al., 2023a). In addition, they fine-tuned PaLM (Chowdhery et al., 2023) to generate both an answer and a URL pointing to Wikipedia page supporting the answer in generative retrieval style (Tay et al., 2022; Wang et al., 2022). Although this setup is similar to ours in that we require the LM to generate the document identifier as well, their setup is basically a variation of RAG where the LM acts as the retriever.
@@ 3 Source-Aware Training
Our training framework is designed to easily integrate with existing pretraining pipelines. We minimize its deviations from established post-pretraining practice, and it involves almost no modifications to the model architecture or implementation. Each document in the pretraining corpus is assigned a unique document identifier (ID) and our goal is to train a language model that can respond to user prompts by providing both a response and an ID referring to the source document of the model’s knowledge.
@@
Our evaluation follows the attributed question answering setup (Bohnet et al., 2022), where given an input prompt $z$, the LLM output will consist of a tuple $\langle r,c\rangle$ where $r$ is the response (e.g., the answer to a question) and $c$ is the identifier of the document in the pretraining data that supports the answer. Following standard LLM training setups, our recipe has two stages: Continual pretraining (Section 3.1) and instruction tuning (Section 3.2). Instruction tuning trains the model to be able to attribute the generated responses to supporting documents it has seen during pretraining. The pretraining stage will involve all documents by nature, but the instruction tuning step is restricted to a subset of the pretraining documents. This restriction is due to the potential cost of curating instruction tuning data from all the pretraining documents, that is in addition to the training overhead incurred by instruction tuning (Zhou et al., 2024).
@@
After training, we measure out-of-domain (OOD) attribution: whether the model can attribute knowledge to documents that are only included in the continual pretraining data but not in the instruction tuning data. We therefore split the pretraining corpus into in-domain and OOD subsets. The in-domain data is used to create attribution training examples, while the OOD documents are used for evaluation, as shown in Figure 2.
3.1   Continual Pretraining with Doc ID Injection
@@
The continual pretraining phase has two goals: (i) memorizing knowledge via next-word prediction (same as established LLM pretraining), and (ii) associating knowledge within a source document with its ID to enable OOD attribution. We aim to achieve the second goal by injecting the document ID into the document before training. An important consideration is the location and frequency of injecting the document ID.
@@
Attention masking to improve training efficiency.
To maximize GPU utilization during continual pretraining, the typical practice packs several pretraining documents within a single training sequence separated by the end-of-sentence <eos> token. As a result, the doc ID tokens for a certain document will naturally attend to preceding tokens from other documents. Our initial experiments showed that this severely hurts attribution, since the model will associate the doc ID of a given document with tokens from other documents in the same training sequence. To avoid this, we modify the causal self-attention mask during pretraining such that the ID tokens for a given document only attend to tokens from within that document.
@@ 3.2   Instruction Tuning
In addition to pretraining, we further adapt the model to (i) recall the appropriate knowledge as a response to the prompt and (ii) cite the ID of the document supporting the response. [^4] The instruction tuning examples are curated from the pretraining data such that for a given prompt, we already have the reference document the model should cite. More details are in Section 4. This stage does not teach the model any new knowledge, but merely aims at eliciting memorization of both knowledge and doc ID by instruction tuning (Wei et al., 2021; Zhang et al., 2023). Given $l$ examples, the $i$-th example is a tuple $\langle z^{(i)},r^{(i)},c^{(i)}\rangle$, where $z^{(i)}$ is the prompt (instruction + query), $r^{(i)}$ is a ground-truth response, and $c^{(i)}$ is the ID of a document that supports the response. The model is trained with the objective $\max_{\theta}\sum_{i=1}^{l}\log P(r^{(i)}|z^{(i)};\theta)P(c^{(i)}|r^{(i)},z^{(i)};\theta)$. The instruction-tuning examples only come from the in-domain documents, and we use the instruction “Answer the following question and provide evidence for your answer.” Figure 2 shows a fine-tuning example from BioCite. During the standard LLM pretraining, i.e., with no-id, we remove the doc ID part from instruction tuning examples. Following Taylor et al. (2022), we surround document IDs with two learned special tokens <id> and </id> during both pretraining and fine-tuning.
@@ 3.3   Chain-of-Thought Attribution
The setup in the previous section train the model to recall the doc ID immediately after the answer. Another setup we explore is where the model is asked to cite the ground-truth document (or part of it) before generating the doc ID. This can be thought of as an instance of chain-of-thought (CoT) (Nye et al., 2021; Wei et al., 2022). In CoT, we inject the document using doc-end, then the model is trained to cite the rest of the document after the answer up till the doc ID. Figure 6 in the Appendix shows an example of the CoT setup.
@@ 4 Data
BioCite is based on the BioS dataset (Zhu & Li, 2023), which is a collection of biographies of fake people where each biography lists six different facts about each person: birthdate, birth city, study major, university, employer, and work city. [^5] Details about reproducing BioS are in Section A.1. Each attribute is described using a corresponding template. For example, the birth city fact is described by “<person name> was born in <birth city>.” To avoid co-reference issues when sampling facts, the person’s full name is mentioned in all the facts.
@@
To simulate realistic pretraining data that often include facts about different entities, we construct each document in BioCite as a collection of facts from at least two different biographies in BioS. More particularly, to construct one document $d$, we first sample the number of biographies $n_{d}\sim\text{Uniform}([2,\cdots,N_{\text{MaxNBio}}])$. Then, we sample $n_{d}$ biographies from BioS without replacement. Finally, we sample a random number of facts from each one in the $n_{d}$ biographies and combine these to form the document. We allow the same combination of biographies to create a document only once and allow each fact to appear only once in BioCite. [^6] In this work, we assume each fact in BioCite is mentioned in exactly one document and leave the extension of this work to multi-doc citation to future work. In our experiments, we generate 100K documents in total using $N_{\text{MaxNBio}}=4$.
@@
It has been shown that the document ID design plays a role in generative retrieval performance (Tay et al., 2022; Pradeep et al., 2023; Sun et al., 2024) and we observed the same during our initial experiments. When designing a doc ID, we need to be careful not to make the task too easy, where the model can infer the doc ID from the input question without actually performing attribution. The design of our dataset allows us to use the last names of the individuals included in a document for two reasons. First, two facts from the same person will most likely exist in many different documents. Second, the same last name can be shared by many different biographies, whose individuals differ only in the first name. That means relying on the last name will not be sufficient to predict the correct doc ID. We choose to use a dash-separated concatenation of the 3-letter prefixes of the last names from the biographies that make up the document, shuffled randomly. We analyze the model predictions when prompted with inputs sharing the same person’s last name in Section 5.3. Table 1 shows an example document, its ID, and a question extracted from it. Exact dataset statistics are in Table 6 in the Appendix.
@@ 5.2   Results
repeat may be unfavorable, since the number of pretraining tokens will noticeably increase by about 80 tokens, bringing additional training overhead. Besides, the model quality will be negatively impacted since document IDs are not natural text, which is reflected in the perplexity over Wikitext-v2 shown in Figure 3 (right). The question here is whether source-aware training can yield OOD attribution while injecting the doc ID once. Interestingly, the chain-of-thought setup (Section 3.3) achieves reasonable OOD attribution without requiring repeating the doc ID within the document. It is worth noting, however, that the CoT setup adds extra training and inference overhead required to generate chain part of the output. Another interesting observation is that repeat and doc-end + CoT achieve better OOD answer EM compared to no-id (e.g., 88.8% with repeat vs. 80.9% with no-id). We conjecture that source-aware training improves the model grounding to the pretraining data, which reflects on the QA performance.
@@ 5.3   Additional Analysis
Now we study the impact of different document ID injection strategies on the LLM quality measured in terms of perplexity over Wikitext-v2. Figure 3 (right) shows perplexity trends during both pretraining and instruction tuning over BioCite and Figure 5 (Left) visualizes the tradeoff between LLM quality and OOD attribution. First, we note that perplexity increases during training in all setups due to the domain shift incurred by training BioCite, which does not resemble real text. We can use the perplexity with no-id as a baseline and observe how other setups compare to it.
Figure 4: The effect of document- and fact-level augmentation on OOD answer match and Hits@$1$. $N_{\text{aug}}=1$ means no augmentation applied. We show results when training with the doc ID injection strategy repeat. Document-level augmentation is necessary for OOD attribution.
@@
Since the doc IDs are constructed as a concatenation of the first three letters of last names in the facts in the documents, the LLM could shortcut the process by predicting doc IDs that contain the prefix of the last name in the question. To verify, we compute the average overlap in the top 10 predicted doc IDs for every pair of OOD questions that share the same last name. We obtain a very low Jaccard Index of 0.08, showing that the model is mostly relying on the whole input rather than only the last name. Table 2 shows two examples of such outputs and the top three predicted doc IDs for each question.
@@
Answer: University of Pittsburgh.
Answer: New Orleans.
@@
Gold document: Adelyn West was born on August 7, 1954. Alissa West lives in New Orleans. Adelyn West studied at University of Pittsburgh…
Gold document: Angelina Grimes was born on December 27, 1916. Angelina Grimes studied at…
Table 2: Two OOD input questions with the same last name and the model predictions. The model predicts totally different doc IDs showing that the model is relying on the fact content itself rather than merely predicting IDs that contain the last name prefix.

### v3

@@ Abstract
Large language models (LLMs) learn a vast amount of knowledge during pretraining, but they are often oblivious to the source(s) of such knowledge. We investigate the problem of intrinsic source citation, where LLMs are required to cite the pretraining source supporting a generated response. Intrinsic source citation can enhance LLM transparency, interpretability, and verifiability. To give LLMs such ability, we explore source-aware training—a recipe that involves (i) training the LLM to associate unique source document identifiers with the knowledge in each document, followed by (ii) an instruction-tuning stage to teach the LLM to cite a supporting pretraining source when prompted. Source-aware training borrows from existing pretraining/fine-tuning frameworks and requires minimal changes to the model architecture or implementation. Through experiments on synthetic data, we demonstrate that our training recipe can enable faithful attribution to the pretraining data without a substantial impact on the model’s perplexity compared to standard pretraining. Our findings also highlight the importance of pretraining data augmentation in achieving attribution. [^1] Code and data available here: https://github.com/mukhal/intrinsic-source-citation.
@@ 1 Introduction
We focus on intrinsic source citation, where the LLM should cite source documents from the pretraining data from which it acquired its relevant parametric knowledge. Compared to retrieval-based approaches such as RAG (Lewis et al., 2020; Guu et al., 2020) or post-hoc techniques (He et al., 2023; Gao et al., 2023a), intrinsic source citation is inherently tied to the model itself, enabling more faithful attribution to its parametric knowledge, and opening up unique opportunities for improved interpretability (Alvarez Melis & Jaakkola, 2018; Marasovic et al., 2022).
To this end, we explore source-aware training to enable an LLM to cite the pretraining source supporting its parametric knowledge. Our motivation is three-fold. First, a significant portion of an LLM’s knowledge is acquired during pretraining, therefore citing evidence for this parametric knowledge can greatly enhance the LLM trustworthiness. Second, the standard practice for LLM pretraining neglects attribution, which explains why LLMs currently fail to provide reliable citations (Agrawal et al., 2023; Zuccon et al., 2023). We aim to explore a training procedure that naturally facilitates citation of the pretraining data. Finally, from a scientific perspective, it is intriguing to investigate whether and how current language models can be trained to reference their pretraining data.
We inquire: Given an off-the-shelf LLM, can we train it to attribute its generations to the supporting sources from the pretraining data? Our goal is to cite the pretraining documents themselves (see Figure 1). Our setup mirrors existing frameworks for LLM (continual) pretraining and can be summarized as follows: We take an off-the-shelf LLM, continue to pretrain it on a corpus associating each document with a unique identifier, then fine-tune it to answer questions about the acquired knowledge while providing citations. The citation is achieved by generating an identifier of a document supporting the answer. Continual pretraining is done as in prior work, with the main difference of injecting the document identifiers into the pretraining data—minimal changes in the model’s architecture or implementation are needed.
To study the generalization over this task and simulate a realistic fine-tuning setting, we limit our instruction tuning stage to a subset of the pretraining documents (in-domain) and evaluate the model’s attribution ability over the remaining (out-of-domain) documents. We run experiments over a synthetic pretraining corpus of fictitious biographies and show that LLMs can achieve reasonable attribution when answering a question about the out-of-domain documents.
@@
  • We explore source-aware training, which we apply to off-the-shelf LLMs to give them the ability to attribute their outputs to the pretraining sources. On synthetic data, we show that such training can achieve reasonable attribution without strongly hurting the LLM perplexity compared to standard pretraining.
  • We examine the impact of various training strategies on attribution, such as training to cite the gold document (Figure 3) and data augmentation (Section 5.3), and our findings can inform future efforts to train attribution-capable models at a large scale.
@@ 2 Related Work
Model-based techniques involve prompting the model directly to generate citations for its parametric knowledge (Weller et al., 2023; Zuccon et al., 2023) or scaling techniques such as influence functions (Koh & Liang, 2017) to large models (Grosse et al., 2023). Model-based attribution is arguably more faithful than retrieval-based approaches as the citation mechanism is intrinsic to the model (Alvarez Melis & Jaakkola, 2018; Marasovic et al., 2022). However, standard approaches to pretraining LMs do not take into account the need for the language model to cite its pretraining data, which is where our work comes into play.
Bohnet et al. (2022) proposed the task of attributed question-answering and evaluated the attribution performance of different systems using the AutoAIS metric (Rashkin et al., 2023; Gao et al., 2023a). In addition, they fine-tuned PaLM (Chowdhery et al., 2023) to generate both an answer and a URL pointing to Wikipedia page supporting the answer in generative retrieval style (Tay et al., 2022; Wang et al., 2022). Although this setup is similar to ours in that we require the LM to generate the document identifier as well, their setup is basically a variation of RAG where the LM acts as the retriever.
@@ 3 Source-Aware Training
Our training framework is designed to easily integrate with existing pretraining pipelines. We minimize its deviations from established pretraining practice, and it involves almost no modifications to the model architecture or implementation. Each document in the pretraining corpus is assigned a unique document identifier (ID) and our goal is to train a language model that can respond to user prompts by providing both a response and an ID referring to the source document of the model’s knowledge.
@@
Our evaluation follows the attributed question answering setup (Bohnet et al., 2022), where given an input prompt $z$, the LLM output will consist of a tuple $\langle r,c\rangle$ where $r$ is the response (e.g., the answer to a question) and $c$ is the identifier of the document in the pretraining data that supports the answer. Following standard LLM training setups, our recipe has two stages: Pretraining (Section 3.1) and instruction tuning (Section 3.2). Instruction tuning trains the model to be able to attribute the generated responses to supporting documents it has seen during pretraining. The pretraining stage will involve all documents by nature, but the instruction tuning step is restricted to a subset of the pretraining documents. This restriction is due to the potential cost of curating instruction tuning data from all the pretraining documents, that is in addition to the training overhead incurred by instruction tuning (Zhou et al., 2024).
@@
After training, we measure out-of-domain (OOD) attribution: whether the model can attribute knowledge to documents that are only included in the pretraining data but not in the instruction tuning data. We therefore split the pretraining corpus into in-domain and OOD subsets. The in-domain data is used to create instruction training examples, while the OOD documents are used for evaluation, as shown in Figure 2.
3.1   Pretraining with Doc ID Injection
@@
The pretraining phase has two goals: (i) memorizing knowledge via next-word prediction (same as established LLM pretraining), and (ii) associating knowledge within a source document with its ID to enable OOD attribution. We aim to achieve the second goal by injecting the document ID into the document before training. An important consideration is the location and frequency of injecting the document ID.
@@
Disabling cross-document attention.
To maximize GPU utilization during continual pretraining, the typical practice packs several pretraining documents within a single training sequence separated by the end-of-sentence <eos> token. As a result, if cross-document attention is enabled with causal masking, the doc ID tokens for a certain document will attend to preceding tokens from other documents and vice versa. Our initial experiments showed that this severely hurts attribution, since the model will associate the doc ID of a given document with knowledge from other documents in the same training sequence. To avoid this, we disable cross-document attention in all our experiments.
@@ 3.2   Instruction Tuning
In addition to pretraining, we further adapt the model to (i) recall the appropriate knowledge as a response to the prompt and (ii) cite the ID of the document supporting the response. [^4] The instruction tuning examples are curated from the pretraining data such that for a given prompt, we already have the reference document the model should cite. More details are in Section 4. This stage does not teach the model any new knowledge, but merely aims at eliciting memorization of both the knowledge and doc ID by instruction tuning (Wei et al., 2021; Zhang et al., 2023). Given $l$ examples, the $i$-th example is a tuple $\langle z^{(i)},r^{(i)},c^{(i)}\rangle$, where $z^{(i)}$ is the prompt (instruction + query), $r^{(i)}$ is a ground-truth response, and $c^{(i)}$ is the ID of a document that supports the response. The model is trained with the objective $\max_{\theta}\sum_{i=1}^{l}\log P(r^{(i)}|z^{(i)};\theta)P(c^{(i)}|r^{(i)},z^{(i)};\theta)$. The instruction-tuning examples only come from the in-domain documents, and we use the simple instruction “Answer the following question and provide evidence for your answer.” Figure 2 shows a fine-tuning example from BioCite. During the standard LLM pretraining, i.e., with no-id, we remove the doc ID part from instruction tuning examples. Following Taylor et al. (2022), we surround document IDs with two learned special tokens <id> and </id> during both pretraining and fine-tuning.
@@ 3.3   Chain-of-Thought Attribution
The setup in the previous section trains the model to recall the doc ID immediately after the answer. Another setup we explore is where the model is asked to cite the ground-truth document (or part of it) before generating the doc ID. This can be regarded as an instance of chain-of-thought (CoT) (Nye et al., 2021; Wei et al., 2022). In CoT, we inject the document using doc-end, then the model is trained to cite the rest of the document after the answer up till the doc ID. Figure 3 shows an example of the CoT setup.
@@ 4 Data
BioCite is based on the BioS dataset (Zhu & Li, 2023), a collection of biographies of fake people where each biography lists six different facts about each person: birthdate, birth city, study major, university, employer, and work city. [^5] Details about reproducing BioS are in Section A.1. Each attribute is described using a corresponding template. For example, the birth city fact is described by “<person name> was born in <birth city>.” To avoid co-reference issues when sampling facts, the person’s full name is mentioned in all the facts.
@@
To simulate realistic pretraining data that often include facts about different entities, we construct each document in BioCite as a collection of facts from at least two different biographies in BioS.
More specifically, to construct one document $d$, we first sample the number of biographies $n_{d}\sim\text{Uniform}([2,\cdots,N_{\text{MaxNBio}}])$. Then, we sample $n_{d}$ biographies from BioS without replacement. Finally, we sample a random number of facts from each one in the $n_{d}$ biographies and combine these to form the document. We allow the same combination of biographies to create a document only once and allow each fact to appear only once in BioCite. [^6] In this work, we assume each fact in BioCite is mentioned in exactly one document and leave the extension of this work to multi-doc citation to future work. In our experiments, we generate 100K documents in total using $N_{\text{MaxNBio}}=4$.
@@
It has been shown that the document ID design plays a role in generative retrieval performance (Tay et al., 2022; Pradeep et al., 2023; Sun et al., 2024) and we observed the same during our initial experiments. When designing a doc ID, we need to be careful not to make the task too easy, where the model can infer the doc ID from the input question without actually performing attribution, but also not too hard, where no semantic overlap exists between the doc ID and the content of the document. Using the last names of the individuals included in a document fits these criteria for two reasons. First, two facts from the same person will most likely exist in different documents. Second, the same last name can be shared by different biographies, whose individuals differ only in the first name. That means relying on the last name will not be sufficient to trivially predict the correct doc ID. We choose to use a dash-separated concatenation of the 3-letter prefixes of the last names from the biographies that make up the document, shuffled randomly. Table 1 shows an example document, its ID, and a question extracted from it. Exact dataset statistics are in Table 7 in the Appendix.
@@ 5.2   Results
repeat may be unfavorable, since the number of pretraining tokens will noticeably increase, bringing additional training overhead. [^8] The number of pretraining tokens increased by about 80% with repeat compared to no-id. Besides, the model quality will be negatively impacted since document IDs are not natural text, which is reflected in the perplexity over Wikitext-v2 shown in Figure 4 (right). The question here is whether source-aware training can yield OOD attribution while injecting the doc ID once. Interestingly, the chain-of-thought setup (Figure 3) achieves reasonable OOD attribution without requiring repeating the doc ID within the document. It is worth noting, however, that the CoT setup adds extra training and inference overhead required to generate chain part of the output. Another interesting observation is that repeat and doc-end + CoT achieve better OOD answer EM compared to no-id (e.g., 88.8% with repeat vs. 80.9% with no-id). We conjecture that source-aware training improves the model grounding to the pretraining data, which reflects on the QA performance.
@@ 5.3   Additional Analysis
Now we study the impact of different document ID injection strategies on the LLM quality measured in terms of perplexity over Wikitext-v2. Figure 4 (right) shows perplexity trends during both pretraining and instruction tuning over BioCite and Figure 5 (Left) visualizes the tradeoff between LLM quality and OOD attribution.
First, we note that perplexity increases during training in all setups due to the domain shift incurred by training BioCite, which does not resemble real text. We can use the perplexity with no-id as a baseline and observe how other setups compare to it.
@@
Figure 6: The effect of document- and fact-level augmentation on OOD answer match and Hits@$1$. $N_{\text{aug}}=1$ means no augmentation applied. We show results when training with the doc ID injection strategy repeat.
@@
Random doc IDs.
EM Hits@1
Last names 88.8 47.3
Random ID 95.9 39.8
Table 2: Out-of-domain QA and attribution performance with different document ID designs.
To study the effect of the doc ID design, we run experiments with randomly generated doc IDs, where each ID is an 8-digit string (e.g., 53476421). We compare the performance of random IDs to our last name-based design when using repeat. Table 2 shows the out-of-domain answer exact match and hits@1 on BioCite. While the model still acheives reasonable attribution with random IDs, they underperform our truncated last name design since it is harder for the model to associate numerical tokens (e.g., the token “3”) than natural language tokens (e.g., “rog”) with document facts.
@@
Since the doc IDs are constructed as a concatenation of the first three letters of last names in the facts in the documents, the LLM could cheat by predicting doc IDs that contain the prefix of the last name in the question. If that were the case, then given two questions about two individuals with the same last name (e.g., Q1: when was Maria Thomson born?; Q2: where does Jake Thomson live?), their generated doc IDs would significantly overlap. We measure this by predicting the top 10 doc IDs for every out-of-domain question. Then for every pair of questions about individuals sharing the same last name (there are 352K such pairs), we compute the overlap between the top 10 predicted doc IDs corresponding to each question using Jaccard index. The average Jaccard over all such pairs is only 0.08, indicating that this is not the case. Table 3 shows two examples of such outputs and the top three predicted doc IDs for each question.
@@
Answer: University of Pittsburgh. Answer: New Orleans.
@@
Gold document: Adelyn West was born on August 7, 1954. Alissa West lives in New Orleans. Adelyn West studied at University of Pittsburgh… Gold document: Angelina Grimes was born on December 27, 1916. Angelina Grimes studied at…
Table 3: Two out-of-domain input questions with the same last name and the model predictions. The model predicts totally different doc IDs showing that the model is relying on the fact content itself rather than merely predicting IDs that contain the last name prefix.